In a biology class, a group of students brought 9 mice for their experiment. They measured the weight and the length of the body of each mouse from the tail to the nose. The findings are recorded below:


Length in cm

13

14

14

11

12

13

12

14

14


Weight in kg

10

10

12

11

16

14

14

17

17


What is the slope of the regression equation? Round your answers to the nearest hundredths.


1
Expert's answer
2022-05-31T11:08:15-0400

In order to compute the regression coefficients, the following table needs to be used:


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
 & X & Y & XY & X^2 & Y^2 \\ \hline
 & 13 & 10 & 130 & 169 & 100 \\
 \hdashline
 & 14 & 10 & 140 & 196 & 100 \\
 \hdashline
 & 14 & 12 & 168 & 196 & 144 \\
 \hdashline
 & 11 & 11 & 121 & 121 & 121 \\
 \hdashline
 & 12 & 16 & 192 & 144 & 256 \\
 \hdashline
 & 13 & 14 & 182 & 169 & 196 \\
 \hdashline
 & 12 & 14 & 168 & 144 & 196 \\
 \hdashline
 & 14 & 17 & 238 & 196 & 289 \\
 \hdashline
 & 14 & 17 & 238 & 196 & 289 \\
 \hdashline
Sum= & 117 & 121 & 1577 & 1531 & 1691 \\
 \hdashline
\end{array}""\bar{X}=\dfrac{1}{n}\sum _{i}X_i=\dfrac{117}{9}=13"




"\bar{Y}=\dfrac{1}{n}\sum _{i}Y_i=\dfrac{121}{9}=13.44"




"SS_{XX}=\sum_iX_i^2-\dfrac{1}{n}(\sum _{i}X_i)^2""=1531-\dfrac{117^2}{9}=10"




"SS_{YY}=\sum_iY_i^2-\dfrac{1}{n}(\sum _{i}Y_i)^2""=1691-\dfrac{121^2}{9}=64.22"




"SS_{XY}=\sum_iX_iY_i-\dfrac{1}{n}(\sum _{i}X_i)(\sum _{i}Y_i)""=1577-\dfrac{117(121)}{9}=4"




"slope=b=\dfrac{SS_{XY}}{SS_{XX}}=\dfrac{4}{10}=0.40"




"a=\dfrac{121}{9}-0.4(13)=8.24"

The regression equation is:


"Y=8.24+0.40X"